OpenStudy (anonymous):

Find the real or imaginary solutions of each equation by factoring. x^3+64

OpenStudy (anonymous):

I need someone to walk me through this.

OpenStudy (campbell_st):

well this is the sum of 2 cubes the general form is \[a^3 + b^3 = (a + b)(a^2 - ab + b^2)\] so you have \[x^3 + 4^3\] when you factor, you get a linear and quadratic factor so you'll need the general quadratic formula for the the imaginary parts of the quadratic hope it helps

OpenStudy (anonymous):

i did that and when i factored it i got (x-4)(x^2 - 4x +16) i wasn't too sure on how to get the actual solution from there.

OpenStudy (campbell_st):

ok... to get solutions you need the factors to equal zero so looking at the linear factor you can use x + 4 = 0 so x = -4 is a solution now the quadratic \[x^2 - 4x + 16 = 0\] you can either use completing the square or the general quadratic formula from here for completing the square, start with \[x^2 - 4x + ? = -16 + ?\] what do you add to both sides of the equation to complete the square..?

OpenStudy (anonymous):

I have it from here. :) thank you. I would rather use the quadratic formula because my calculator helps with that. thank you for your help.